### v1

# Regularization by regular noise: a numerical result

Abstract.
We study a singular stochastic equation driven by a regular noise of fractional Brownian type with Hurst index $H\in(1,\infty)\setminus\mathbb{Z}$ and drift coefficient $b\in\mathcal{C}^{\alpha}$, where $\alpha>1-\tfrac{1}{2H}$. The strong well-posedness of this equation was first established in [Ger23], a phenomenon referred to as regularization by regular noise. In this note, we provide a corresponding numerical analysis. Specifically, we show that the Euler–Maruyama approximation $X^{n}$ converges strongly to the unique solution $X$ with rate $n^{-1}$. Furthermore, under the additional assumption $b\in\mathcal{C}^{1}$, we show that $n(X-X^{n})$ converges to a non-trivial limit as $n\to\infty$, thereby confirming that the rate $n^{-1}$ is in fact optimal upper bound for this scheme.
Mathematics Subject Classification (2020): Primary 60H35, 65C30, 60H10; Secondary 60H50, 60L90.
Keywords: Singular SDEs; Strong approximation; Euler scheme; Regularization by noise; Stochastic sewing, Fractional Brownian motion.
1. Introduction
It is known that when $H\in(0,1)$, a $d$-dimensional fractional Brwonian motion (fBM) $B^{H}$ can be defined via the Mandelbrot - van Ness representation [MVN68]:
$B_{t}^{H}:=\int_{-\infty}^{0}(|t-s|^{H-\frac{1}{2}}-|s|^{H-\frac{1}{2}})\mathrm{d}W_{s}+\int_{0}^{t}|t-s|^{H-\frac{1}{2}}\mathrm{d}W_{s},\quad t\geqslant 0,$ (1.1)
[…]
$B_{t}^{H}:=\int_{0\leqslant s_{1}\leqslant\ldots\leqslant s_{\lfloor H\rfloor}\leqslant t}B^{H-\lfloor H\rfloor}_{s_{1}}\mathrm{d}{s_{1}}\ldots\mathrm{d}s_{\lfloor H\rfloor}.$ (1.2)
Due to the possible multiple times of integral in (1.2) we can see clearly that the path of $B^{H}$ for $H>1$ is regular, at least $C^{1}$. This fact yields one of the appealing points of [Ger23] which shows the strong well-posedness of following equation with singular $b$:
$\mathrm{d}X_{t}=b(X_{t})\mathrm{d}t+\mathrm{d}B_{t}^{H},\quad X_{0}=x_{0}\in{\mathbb{R}}^{d},$ (1.3)
[…]
$\left\{\begin{array}[]{cc}\mathrm{d}X_{t}&=\big(b(X_{t})+V_{t}^{\lfloor H\rfloor}\big)\mathrm{d}t,\\ \mathrm{d}V_{t}^{\lfloor H\rfloor}&=V_{t}^{\lfloor H\rfloor-1}\mathrm{d}t,\\ \qquad\qquad\qquad\qquad\ldots,&\\ \mathrm{d}V_{t}^{1}&=\mathrm{d}B_{t}^{H-\lfloor H\rfloor},\end{array}\right$. (1.8)
which shares the similar framework from [CdRHM22, HZZZ24, HLL24, Lin24] but with non-Markvin noise when $H\neq k+\frac{1}{2}$ for any $k\in{\mathbb{N}}$.
In this note our motivation is to provide a numerical result of it by considering its Euler-Maruyama (EM) scheme
$\mathrm{d}X_{t}^{n}=b(X_{k_{n}(t)}^{n})\mathrm{d}t+\mathrm{d}B_{t}^{H},\qquad X_{0}^{n}=x_{0}^{n}\in{\mathbb{R}}^{d}$ (1.9)
with $k_{n}(t)\coloneqq\frac{\lfloor nt\rfloor}{n}$.
When $H\in(0,1)$, [BDG21] has shown the strong convergence rate $n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}$ of the scheme (1.9) to (1.3), inside which the Girsanov Theroem and Stochastic Sewing Lemma (SSL) [Lê20] play one of the most important roles. Precisely the methodology inside [BDG21] can be roughly summarized as (take $x_{0}=x_{0}^{n}$)
$\|\sup_{t\in[0,1]}|X_{t}-X^{n}_{t}|\|_{L^{p}_{\omega}}\overset{\text{Girsanov}}{\lesssim}\big\|\int_{0}^{1}b(B_{s}^{H})-b(B_{k_{n}(s)}^{H})\mathrm{d}r\big\|_{L^{p}_{\omega}}\overset{\text{SSL}}{\lesssim}n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}.$
When $H>1$, on one hand, as indicated in [Ger23, GG25] the Girsanov Theorem becomes less helpful; meanwhile PDE tools clearly would not work. Therefore the challenge for showing the convergence of (1.9) to (1.3) lies in two folds comparing with known results: absence of Girsanov and non-Markovianity. Alternatively, although [Ger23] studies the well-posedness only, actually it already hints the possible way to get rid of Girsanov; concerning numerical approximation [BDG24] provides the similar evidence, however it considers the singular SDE driven by an $\alpha$ stable process which is a Markov process. Nevertheless we are able to show convergence rate $n^{-1}$ which is comparable with [BDG21] since $H>1$ here implies $\alpha>1-\frac{1}{2H}>\frac{1}{2}$.
For such obtained convergence rate, naturally we wonder how far it is away from being optimal. Such question on the optimality of EM scheme has been addressed and studied for instance in [MGY23, EMGY25, DGL23a] concerning Brownian noise, [PSS24] for Lévy processes with jumps and [Neu06, NN07, LLT25, HLN16] for fBM with $H<1$. In particular [Neu06] confirms for $H\in(\frac{1}{2},1)$, order $n^{H}$ reaches its optimality of EM scheme (1.9) for $b\in{\mathcal{C}}^{2}$ (twice differentiable) with possible linear growth. Here we complete this result with rate $n^{-1}$ for $H>1$ and less regular $b$ (actually only ${\mathcal{C}}^{1}$). The idea of the proof is straightforward: we show that for $b\in{\mathcal{C}}^{1}$, the following approximation reaches a possibly non-zero limit:
$n(X_{t}-X^{n}_{t})\overset{n\rightarrow\infty}{\rightarrow}c(t)\neq 0,\quad a.s.$
which indicates that for (1.9) the best convergence rate we can expect should be not faster than order $n$. Evidently it also verifies that the rate we obtained actually is optimal.
Lastly, let us mention a few related works that share a similar interest and spirit in the study of numerical approximations. For equations of the type (1.3) with singular drift, convergence results have been established in [NS21, DG20, JM21, BW25] for additive Brownian motion, in [DGL23b, BDG21, GLL25, LL25] for multiplicative Brownian noise, in [BDG24, BWWZ25] for Lévy processes with jumps, and in [BDG21, GHR25] for fractional Brownian motion. A slightly different notion of singularity-namely, piecewise Lipschitz coefficients-has also been investigated, with convergence results obtained in works such as [LS16, MGY20, MGRY25, PSS24]. We must acknowledge that this is only a brief selection from a vast body of literature, and we encourage readers to explore the cited works for further details and insights.
Organization of the paper
In Section˜2 we introduce the necessary notations and main results. Section˜3 collects all of the crucial tools and properties of fBM. We present the central analysis and proof of the main result in Section˜4. In the end we discuss optimality in Section˜5.
2. Preliminaries and main results
2.1. Preliminaries
On finite dimensional vector spaces we always use the Euclidean norm.
For $k\in\mathbb{N}$, $f:\mathbb{R}^{d}\mapsto\mathbb{R}$, denote $\partial_{k}f(x)\coloneqq\frac{\partial f(x)}{\partial x_{k}}$ for $x\in\mathbb{R}^{d}$ and $\nabla f(x)\coloneqq(\partial_{i}f(x))_{1\leqslant i\leqslant d}$, the derivative is understood in the weak sense. For vector-valued $f$ we use the same notation, and $\nabla^{k}f$ is defined via $\nabla(\nabla^{k-1}f)$ iteratively. For a multi-index $k=(k_{1},\ldots,k_{d})\in{\mathbb{N}}^{d}$, denote $\partial^{k}f(x)\coloneqq\frac{\partial^{|k|}f(x)}{\partial x_{k_{1}}\cdots\partial x_{k_{d}}}$. If $k=(0,\ldots,0)$, we use convention $\partial^{k}f=f$. We denote by ${C}^{\infty}_{0}$ (${C}_{p}^{\infty}$, resp.) the set of all continuously infinitely differentiable functions that, along with all of their partial derivatives, are compactly supported (of polynomial growth, resp.).
For $\alpha\in(0,1]$, we set $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ to be the space of continuous functions such that
$\|f\|_{\mathcal{C}^{\alpha}}\coloneqq[f]_{\mathcal{C}^{\alpha}}+\sup_{x\in{\mathbb{R}}^{d}}|f(x)|\coloneqq\sup_{x,y\in{\mathbb{R}}^{d},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in{\mathbb{R}}^{d}}|f(x)|<\infty.$
Here, and often below, we write $\mathcal{C}^{\alpha}$ instead of $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ for simplicity. For $\alpha\in(0,\infty)$, we define $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ the space of all functions $f$ defined on ${\mathbb{R}}^{d}$ having bounded derivatives $\partial^{k}f$ for multi-indices $k\in{\mathbb{N}}^{d}$ with $|k|\leqslant\alpha$ so that
$\|f\|_{\mathcal{C}^{\alpha}}$ $\coloneqq\|f\|_{\mathcal{C}^{\lfloor\alpha\rfloor}}+[f]_{\mathcal{C}^{\alpha}}\coloneqq\sum_{|k|\leqslant\lfloor\alpha\rfloor}\sup_{x\in{\mathbb{R}}^{d}}|\partial^{k}f(x)|+\sum_{k=\lfloor\alpha\rfloor}[\partial^{k}f]_{\mathcal{C}^{\{\alpha\}}}<\infty,$
where $\{\alpha\}:=\alpha-\lfloor\alpha\rfloor$. Note that the $\mathcal{C}^{\alpha}$-norm always includes the supremum of the function. We also denote the space of bounded measurable functions $\mathcal{C}^{0}({\mathbb{R}}^{d})$ with the supremum norm. To be noticed that the functions in $\mathcal{C}^{0}$ do not need to be continuous.
In the following we denote the conditional expectation w.r.t. the $\sigma$-algebras of the filtration $(\mathcal{F}_{t})_{t\geqslant 0}$ as ${\mathbb{E}}^{t}(\cdot)\coloneqq{\mathbb{E}}(\cdot|\mathcal{F}_{t}),t\geqslant 0$, $\|X\|_{L^{p}_{\omega}}:=({\mathbb{E}}|X|^{p})^{\frac{1}{p}}$, $\|X\|_{L^{p}_{\omega}|{\mathcal{F}}_{s}}:=({\mathbb{E}}[|X|^{p}|{\mathcal{F}}_{s}])^{\frac{1}{p}}$.
For $p\in[1,\infty]$, $X\in L^{p}(\Omega,{\mathbb{R}}^{d})$ and $\mathcal{F}_{s}$-measurable ${\mathbb{R}}^{d}$ valued random vector $Y$, we have the following inequalities
$\|{\mathbb{E}}^{s}X\|_{L^{p}_{\omega}}\leqslant\|X\|_{L^{p}_{\omega}}$ (2.1)
and
$\|X-{\mathbb{E}}^{s}X\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant 2\|X-Y\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\quad a.s.$ (2.2)
Let $f:[0,1]\times\Omega\rightarrow\mathbb{R}^{d}$ be a measurable function adapted to the filtration $(\mathcal{F}_{t})_{t\geqslant 0}$, $\gamma\in(0,1]$, $p\geqslant 2$ and $S,T]\subset[0,1]$. We give the following definitions:
$\|f\|_{C_{p}^{0}[S,T]}\coloneqq\sup_{r\in[S,T]}\|f(r)\|_{L^{p}_{\omega}};$
$[f]_{C_{p}^{\gamma}[S,T]}\coloneqq\sup_{r_{1},r_{2}\in[S,T],r_{1}\neq r_{2}}\frac{\|\partial^{\lfloor\gamma\rfloor}f(r_{1})-\partial^{\lfloor\gamma\rfloor}f(r_{2})\|_{L^{p}_{\omega}}}{|r_{1}-r_{2}|^{\{\gamma\}}};$
$\|f\|_{C_{p}^{\gamma}[S,T]}\coloneqq\|f\|_{C_{p}^{0}[S,T]}+[f]_{C_{p}^{\gamma}[S,T]}.$
If $f$ is an adapted process, we choose $Y$ in (2.2) as the value at $t$ of the Taylor expansion of $f$ at $s$ up to order $\lfloor\gamma\rfloor$ and we obtain
$\|f_{t}-{\mathbb{E}}^{s}f_{t}\|_{L^{p}_{\omega}}\leqslant 2|t-s|^{\gamma}[f]_{{C}^{\gamma}_{p}[s,t]}.$ (2.3)
In proofs, the notation $a\lesssim b$ (respectively $a\asymp b$) abbreviates the existence of $C>0$ such that $a\leqslant Cb$ (respectively $C^{-1}b\leqslant a\leqslant Cb$), such that moreover $C$ depends only on the parameters claimed in the corresponding statement.
2.2. Main results
Our main assumption and result can be stated as follows.
Assumption 2.1.
[…]
Notice that following [Ger23], under the assumption above, there exists a unique strong solution to (1.3). Here is our numerical approximation result for it.
Theorem 2.2.
Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly. Suppose Assumption˜2.1 holds. Then for every $p>1$, we have
$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[0,1]}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1},$ (2.4)
where $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Remark 2.3.
As an easy application of Kolmogorov continuity criteria we can also conclude from (2.4) the following:
$\big\|\sup_{t\in[0,1]}|X_{t}-X_{t}^{n}|\big\|_{L_{\omega}^{p}}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1}.$
Idea of the analysis. Its detailed proof will be given in Section˜4. Here we only outline the rough idea.
Observe that for any $p\geqslant 1$, for $0\leqslant s\leqslant t\leqslant 1$
$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})|\big\|_{L^{p}_{\omega}}=\big\|(\varphi_{t}-\varphi_{t}^{n})-(\varphi_{s}-\varphi_{s}^{n})\big\|_{L^{p}_{\omega}},$
where
$\varphi_{t}:=X_{t}-B_{t}^{H}=\int_{0}^{t}b(X_{r})\mathrm{d}r=\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
$\varphi_{t}^{n}:=X_{t}^{n}-B_{t}^{H}=\int_{0}^{t}b(X_{k_{n}(r)}^{n})\mathrm{d}r=\int_{0}^{t}b(\varphi_{k_{n}(r)}^{n}+B_{k_{n}(r)}^{H})\mathrm{d}r,$
meanwhile, we know that our aim is to get rid of ${\mathcal{C}}^{1}$-regularity demand on $b$ via proper implement of the regularization from $B^{H}$, precisely the Gaussian density $p_{c(H)t^{2H}}$ which is infinitely many smoothing (see Section˜3.2). Also keep in mind in the current setting Girsanov Theorem is not available. What we have learned from [Ger23, BDG24] is that we can achieve this goal via freezing the exponent $\varphi_{r}$ inside the integral $\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ (the same for $\varphi^{n}_{k_{n}(r)}$ inside the integral $\int_{0}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$) via taking conditional expectation in the framework of SSL. That is to say, heuristically for tst-s being small enough
$\int_{s}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
$\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$ $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r,$
and this $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}$” is verified by SSL (see Lemma˜3.1 below) by taking
$A_{s,t}:=\int_{s}^{t}{\mathbb{E}}^{s-(t-s)}b({\mathbb{E}}^{s-(t-s)}\varphi_{r}+B_{r}^{H})\mathrm{d}r,\quad{\mathcal{A}}_{s,t}:=\int_{s}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r,$
similarly for $\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$, then together with the property of Gaussian density $p_{c(H)t^{2H}}$ of fBM ( $\mathcal{P}_{t}^{H}f:=p_{c(H)t^{2H}}\ast f$) we can further write
$A_{s,t}=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})\mathrm{d}r.$
Now we can see that instead of dealing with $b$ directly, we have more regularity to spare in $\mathcal{P}^{H}_{t}b$ due to the smoothing effect from the convolution with $p_{c(H)t^{2H}}$.
Although the true analysis later also contains large part technical settlement, the core of the whole argument is clear. In the end we are able to adjust the idea above into the proof of the convergence rate in the following form:
$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|\varphi_{t}-\varphi_{t}^{n}\big\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
for sufficiently small $\varepsilon$. Therefore we get (2.4) after fine dividing the interval $[0,1]$ and for each sub-interval applying what we have obtained above.
3. Auxiliary Tools
In this section, we primarily introduce our main tool which is the stochastic sewing lemma and present some properties of the fractional Brownian motions.
3.1. Stochastic Sewing Lemma
Given $M\geqslant 0$ we define $S,T]_{M}^{2}=\{(s,t)|S\leqslant s<t\leqslant T,s-M(t-s)\geqslant S\}$ and $\overline{[S,T]}_{M}^{3}=\{(s,u,t)|(s,t)\in[S,T]_{M}^{2},(u-s)\wedge(t-u)\geqslant\frac{t-s}{3}\}$.
Lemma 3.1.
[Ger23, Lemma 2.2] Let $0\leqslant S<T\leqslant 1,p\in[2,\infty),M\geqslant 0$ and let $\left(A_{s,t}\right)_{(s,t)\in[S,T]_{M}^{2}}$ be a family of random variables in $L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $A_{s,t}$ is $\mathcal{F}_{t}$-measurable. Suppose that for some $\varepsilon_{1},\varepsilon_{2}>0$ and $C_{1},C_{2}$ the bounds
$\big\|A_{s,t}\big\|_{L^{p}_{\omega}}\leqslant C_{1}|t-s|^{1/2+\varepsilon_{1}}$ (3.1)
and
$\big\|\mathbf{E}^{s-M(t-s)}\delta A_{s,u,t}\big\|_{L^{p}_{\omega}}\leqslant C_{2}|t-s|^{1+\varepsilon_{2}}$ (3.2)
hold for all $(s,t)\in[S,T]_{M}^{2}$ and $(s,u,t)\in\overline{[S,T]}_{M}^{3}$, where $\delta A_{s,u,t}:=A_{s,t}-A_{s,u}-A_{u,t}$. Then there exists a unique (up to modification) adapted process $\mathcal{A}:[S,T]\rightarrow L^{p}\left(\Omega,\mathbb{R}^{d}\right)$ such that $\mathcal{A}_{S}=0$ and such that for some constants $K_{1},K_{2}<\infty$, depending only on $\varepsilon_{1},\varepsilon_{2},p,d$, and $M$, the bound
$\left\|\mathcal{A}_{t}-\mathcal{A}_{s}\right\|_{L^{p}_{\omega}}\leqslant K_{1}C_{1}|t-s|^{1/2+\varepsilon_{1}}+K_{2}C_{2}|t-s|^{1+\varepsilon_{2}}$ (3.3)
holds for all $(s,t)\in[S,T]_{0}^{2}$. Moreover, if there exists any continuous process $\widetilde{\mathcal{A}}$ : $S,T]\rightarrow L^{p}\left(\Omega,\mathbb{R}^{d}\right),\varepsilon_{3}>0$, and $K_{3}<\infty$, such that $\widetilde{\mathcal{A}}_{S}=0$ and
$\big\|\widetilde{\mathcal{A}}_{t}-\widetilde{\mathcal{A}}_{s}-A_{s,t}\big\|_{L^{p}_{\omega}}\leqslant K_{3}|t-s|^{1+\varepsilon_{3}}$ (3.4)
holds for all $(s,t)\in[S,T]_{M}^{2}$, then $\widetilde{\mathcal{A}}_{t}=\mathcal{A}_{t}$ for all $S\leqslant t\leqslant T$.
3.2. Fractional Brownian motions
Here we recall the following properties concerning $B^{H}$ that have been used heavily in later analysis.
Lemma 3.2.
[…]
${\mathbb{E}}|B_{t}^{H}-{\mathbb{E}}^{s}B_{t}^{H}|^{2}=dc(H)|t-s|^{2H}\text{ and }B_{t}^{H}-{\mathbb{E}}^{s}B_{t}^{H}\text{ is independent of }\mathcal{F}_{s}.$ (3.5)
[…]
${\mathbb{E}}|B_{t}^{H}-B_{s}^{H}|\leqslant N|t-s|^{H\wedge 1}.$ (3.6)
We let $p_{t}(x)$ denote the known heat density $\frac{1}{(2\pi t)^{d/2}}e^{-\frac{|x|^{2}}{2t}}$ on ${\mathbb{R}}^{d}$ and we define ${\mathcal{P}}_{t}^{H}f(x):=(p_{c(H)t^{2H}}\ast f)(x)$, $x\in{\mathbb{R}}^{d}$. Then for any $\mathcal{F}_{s}$-measurable ${\mathbb{R}}^{d}$ valued random vector $\xi$, we have
${\mathbb{E}}^{s}f(B_{t}^{H}+\xi)={\mathcal{P}}_{t-s}^{H}f({\mathbb{E}}^{s}B_{t}^{H}+\xi).$ (3.7)
Lemma 3.3.
For $\alpha,\beta\in[0,1],f\in{\mathcal{C}}^{\alpha},t\in(0,1],$ one has the bounds, with some constant $N$ depending only on $H,\alpha,\beta,d$, accordingly
$|\mathcal{P}_{t}^{H}f(x_{1})-\mathcal{P}_{t}^{H}f(x_{2})-\mathcal{P}_{t}^{H}f(x_{3})+\mathcal{P}_{t}^{H}f(x_{4})|$
$\leqslant N\|f\|_{{\mathcal{C}}^{\alpha}}\big(t^{H(\alpha-2)}|x_{1}-x_{2}||x_{1}-x_{3}|+t^{H(\alpha-1)}|x_{1}-x_{2}-x_{3}+x_{4}|\big),\forall x_{i}\in{\mathbb{R}}^{d},i=1,\ldots,4;$ (3.8)
$\|\mathcal{P}_{t}^{H}f\|_{{\mathcal{C}}^{\beta}}\leqslant Nt^{H(\alpha-\beta)\wedge 0}\|f\|_{{\mathcal{C}}^{\alpha}};$ (3.9)
$\|(\mathcal{P}_{t}^{H}-\mathcal{P}_{s}^{H})f\|_{{\mathcal{C}}^{\beta}}\leqslant Ns^{H(\alpha-\beta)-2H\delta}|t^{2H}-s^{2H}|^{\delta}\|f\|_{{\mathcal{C}}^{\alpha}},\forall 0\leqslant s\leqslant t\leqslant 1,0<\delta\in\left[\frac{\alpha-\beta}{2},1\right].$ (3.10)
Proof.
(3.3) are directly from [Ger23, Page 2 (2.8)]. By using properties of Gaussian convolutions, heat kernel bounds and a relation of the form ${\mathcal{P}}_{t}^{H}f(x)=(p_{c(H)t^{2H}}\ast f)(x)$, we get (3.9) from [BDG21, Proposition 3.7 (i)]. For (3.10), it holds from [BDG21, Proposition 3.7 (ii)]. ∎
4. Strong convergence rate
In this part we give the proof for Theorem˜2.2.
Denote
$\varphi_{t}\coloneqq(X-B^{H})_{t}$ $=x_{0}+\int_{0}^{t}b(\varphi_{s}+B^{H}_{s})\mathrm{d}s,$
$\varphi^{n}_{t}\coloneqq(X^{n}-B^{H})_{t}$ $=x_{0}^{n}+\int_{0}^{t}b(\varphi_{k_{n}(s)}^{n}+B^{H}_{k_{n}(s)})\mathrm{d}s.$
Fix $S\leqslant s<t\leqslant T$ and $S,T]\subset[0,1]$. We write
$(X$ $-X^{n})_{t}-(X-X^{n})_{s}$
$=(\varphi-\varphi^{n})_{t}-(\varphi-\varphi^{n})_{s}$
$=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r})-b(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r+\int_{s}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-b(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$\quad+\int_{s}^{t}b(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$\eqqcolon\mathcal{E}^{b,n,1}_{s,t}+\mathcal{E}^{b,n,2}_{s,t}+\mathcal{E}^{b,n,3}_{s,t}.$
It is clear that in order to show (2.4), we need to estimate ${\mathcal{E}}^{b,n,1}_{s,t},{\mathcal{E}}^{b,n,2}_{s,t},{\mathcal{E}}^{b,n,3}_{s,t}$ individually. We distribute the estimates for each into Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4 correspondingly.
Before that we first present the following auxiliary lemma for the processes $\varphi$ and $\varphi^{n}$ defined above which will be heavily used in the later proofs.
Lemma 4.1.
Assume Assumption˜2.1 holds. Then for all $t>s$ and $p\geqslant 1$ we have
$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.;$ (4.1)
$\|\varphi^{n}_{t}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant N|t-s|^{1+\alpha H}\quad a.s.$ (4.2)
with some constant $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
Suppose (4.1) holds for some $m\geqslant 0$ in place of $1+\alpha H.$ This is certainly true for $m=0$ thanks to the fact that $b$ is bounded; we proceed now by induction on $m.$
We apply (2.2) with $X=\varphi_{t},Y=\varphi_{s}+\int_{s}^{t}b({\mathbb{E}}^{s}B^{H}_{r}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r$ to get
$\|\varphi_{t}-{\mathbb{E}}^{s}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant 2\left\|\varphi_{t}-\varphi_{s}-\int_{s}^{t}b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$=2\left\|\int_{s}^{t}(b(B^{H}_{r}+\varphi_{r})-b({\mathbb{E}}^{s}B_{r}^{H}+{\mathbb{E}}^{s}\varphi_{r}))\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\left\|\int_{s}^{t}(|B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}+|\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(\big\||B^{H}_{r}-{\mathbb{E}}^{s}B^{H}_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}}+\big\||\varphi_{r}-{\mathbb{E}}^{s}\varphi_{r}|^{\alpha}\big\|_{L^{p}_{\omega}|\mathcal{F}_{s}})\mathrm{d}r.$
Using (3.7) and the induction hypothesis, we get
$\|\varphi_{t}-{\mathbb{E}}\varphi_{t}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}\leqslant N|t-s|^{(H\alpha)\wedge(m\alpha)+1}\quad a.s.$
We note that $m_{0}=0,m_{i+1}=1+(H\alpha)\wedge(m_{i}\alpha)$ reaches $1+H\alpha$ in finitely many steps, therefore we get (4.1).
Inequality (4.2) is obtained by a similar argument. For fixed $s,$ define $s^{\prime}$ to be the smallest grid point which is bigger or equal to $s,$ that is, $s^{\prime}:=\lceil ns\rceil n^{-1}$. It is crucial to note that $\varphi_{s^{\prime}}^{n}$ is $\mathcal{F}_{s}$-measurable. We proceed by induction as before and assume that (4.2) holds for some $m\geqslant 0.$ If $s\leqslant t<s^{\prime},$ then $\varphi_{t}^{n}$ is $\mathcal{F}_{s}$-measurable. Hence $\varphi^{n}_{t}={\mathbb{E}}^{s}\varphi_{t}^{n}$ and the left-hand side of (4.2) is zero. Therefore it remains to consider the case $t\geqslant s^{\prime}.$ In this case, using again (2.2) with $X=\varphi_{t}^{n},Y=\varphi^{n}_{s^{\prime}}+\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$ we deduce
$\|\varphi_{t}^{n}-{\mathbb{E}}^{s}\varphi_{t}^{n}\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$ $\leqslant 2\left\|\varphi_{t}^{n}-\varphi^{n}_{s^{\prime}}-\int_{s^{\prime}}^{t}b({\mathbb{E}}^{s}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$=2\left\|\int_{s^{\prime}}^{t}(b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})-b({\mathbb{E}}^{s}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s}\varphi_{k_{n}(r)}^{n})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}$
$\leqslant N\left\|\int_{s^{\prime}}^{t}(|B^{H}_{k_{n}(r)}-{\mathbb{E}}^{s}B^{H}_{k_{n}(r)}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s}\varphi^{n}_{k_{n}(r)}|^{\alpha})\mathrm{d}r\right\|_{L^{p}_{\omega}|\mathcal{F}_{s}}.$
From now on we obtain (4.2) as before. ∎
Here we introduce some notations commonly used in the proofs of Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4. For $(s,u,t)\in\overline{[0,1]}_{1}^{3}$, we set
$s_{1}\coloneqq s-(t-s),s_{2}\coloneqq u-(t-u),s_{3}\coloneqq s-(u-s),s_{4}\coloneqq s,s_{5}\coloneqq u,s_{6}\coloneqq t.$ (4.3)
Note by the fact $u\leqslant\frac{2}{3}s+\frac{1}{3}t$ for $(s,u,t)\in\overline{[0,1]}_{1}^{3}$, we have $s_{2}\leqslant s_{3}.$
Let us start with the estimate for $\mathcal{E}^{b,n,1}$.
Lemma 4.2.
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ we have
$\left\|\mathcal{E}^{b,n,1}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant N\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}|t-s|^{\frac{1}{2}+\varepsilon},\quad\forall(s,t)\in[0,1]_{0}^{2}$ (4.4)
with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
The idea is to apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})-b(B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})\mathrm{d}r.$
We are going to verify (3.1) and (3.2). By (3.7), we see
$A_{s,t}$ $=\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{r})\mathrm{d}r.$ (4.5)
Then by (3.9) and (2.1), we get
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-[s-(t-s)])^{-(1-\alpha)H}\|{\mathbb{E}}^{s-(t-s)}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\mathrm{d}r$ (4.6)
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}|t-s|^{1-(1-\alpha)H}.$
Then (3.1) holds with $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$ by the fact that $1-(1-\alpha)H>\frac{1}{2}$.
Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). We first can write
${\mathbb{E}}^{s-(t-s)}\delta A_{s,u,t}$ (4.7)
$=$ ${\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{3}}\int_{s_{4}}^{s_{5}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+b(B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}{\mathbb{E}}^{s_{2}}\int_{s_{5}}^{s_{6}}b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-b(B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+b(B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
$=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})\mathrm{d}r$
$=:$ $I_{1}+I_{2}.$
The two terms are treated in the exactly same way, so we only detail $I_{1}$. By (2.1) and applying (3.3) with
$x_{1}={\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n},x_{2}={\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r},x_{3}={\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n},x_{4}={\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{r},$
we obtain
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}|r-s_{3}|^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ (4.8)
$\quad\qquad\qquad\quad+|r-s_{3}|^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\mathrm{d}r.$
By (4.2),
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|={\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}(\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})|\leqslant{\mathbb{E}}^{s_{1}}|\varphi^{n}_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant N|r-s_{1}|^{1+\alpha H}.$ (4.9)
Besides from (2.1), we get
$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}$ $=\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}((\varphi_{r}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n}))|\big\|_{L^{p}_{\omega}}$ (4.10)
$\leqslant\big\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})|\big\|_{L^{p}_{\omega}}$
$\leqslant\|(\varphi_{r}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}},$
meanwhile (2.3) implies
$\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}-{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\big\|_{L^{p}_{\omega}}\leqslant N|r-s_{1}|^{\frac{1}{2}}[\varphi-\varphi^{n}]_{C^{\frac{1}{2}}_{p}},$ (4.11)
clearly from (2.1),
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}-\varphi_{r}^{n})\|_{L^{p}_{\omega}}\leqslant\|\varphi-\varphi^{n}\|_{C_{p}^{0}},$ (4.12)
now plugging (4.9), (4.11) and (4.12) into (4.8), we have
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{\frac{1}{2}}(r-s_{3})^{-(1-\alpha)H}\mathrm{d}r$ (4.13)
$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}\int_{s_{4}}^{s_{5}}(r-s_{1})^{1+\alpha H}(r-s_{3})^{-(2-\alpha)H}\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\frac{3}{2}-(1-\alpha)H}[\varphi-\varphi^{n}]_{C_{p}^{\frac{1}{2}}}$
$\quad+\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{2+\alpha H-(2-\alpha)H}\|\varphi-\varphi^{n}\|_{C_{p}^{0}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}(t-s)^{(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)}.$
The above analysis also implies the same bound on $I_{2}$ observing $I_{1}$ and $I_{2}$ share the same structure.
Noticing Assumption˜2.1 implies $(\frac{3}{2}-(1-\alpha)H)\wedge(2+\alpha H-(2-\alpha)H)>1$, we conclude (3.2) holds with $C_{2}=N\|b\|_{{\mathcal{C}}^{\alpha}}\|\varphi-\varphi^{n}\|_{C_{p}^{\frac{1}{2}}}$.
Now we claim that the process $\mathcal{A}$ in (3.3) actually is given by
${\mathcal{A}}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r})-b(B^{H}_{r}+\varphi_{r}^{n})\mathrm{d}r.$ (4.14)
To prove this, it suffices to show (3.4). By (4), we write
${\mathcal{A}}_{t}-{\mathcal{A}}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r$ (4.15)
$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})\mathrm{d}r$
$=:II_{1}+II_{2}.$
Again we can see that $II_{2}$ can be treated similarly to $II_{1}$, so we only detail $II_{1}$. We can see
$II_{1}$ $=\int_{s}^{t}(b-\mathcal{P}^{H}_{r-s_{1}}b)(B_{r}^{H}+\varphi_{r})\mathrm{d}r+\int_{s}^{t}\mathcal{P}^{H}_{r-s_{1}}b(B_{r}^{H}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r.$ (4.16)
Using (3.10) with $\delta=\frac{\alpha}{2}$, $\beta=0$ and (3.9) with $\beta=\alpha$, we get
$\|II_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\int_{s}^{t}\|\mathcal{P}^{H}_{r-s_{1}}b-b\|_{{\mathcal{C}}^{0}}+\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{\alpha}}(\|B_{r}^{H}-{\mathbb{E}}^{s_{1}}B^{H}_{r}\|^{\alpha}_{L^{\alpha p}_{\omega}}+\|\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}\|^{\alpha}_{L_{\omega}^{\alpha p}})\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-s_{1})^{\alpha H}+(|r-s_{1}|^{\alpha H}+|r-s_{1}|^{\alpha(1+\alpha H)})\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{(1+\alpha H)\wedge(1+\alpha(1+\alpha H))},$ (4.17)
where in the second inequality we used (3.6) and (4.1). The same bound on $II_{2}$.
Therefore, (3.4) holds since $(1+\alpha H)\wedge(1+\alpha(1+\alpha H))>1$. Then the uniqueness from Lemma˜3.1 verifies the claim.
Finally, by Lemma˜3.1, the proof is complete. ∎
Let us move to estimate $\mathcal{E}^{b,n,2}$ term.
Lemma 4.3.
Suppose Assumption˜2.1 holds. Then we have for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$
$\left\|\mathcal{E}^{b,n,2}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.18)
where $\epsilon>0$ is sufficiently small and $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
Again the idea is to apply Lemma˜3.1. Let $M=1$, $(s,t)\in[0,1]_{1}^{2}$ and
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.19)
Similarly to (4) and (4.6), we have
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $=\|\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)})\mathrm{d}r\|_{L^{p}_{\omega}}$ (4.20)
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
$\leqslant Nn^{-1}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (4.21)
where in the second inequality we used $\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}\leqslant\|b\|_{{\mathcal{C}}^{0}}n^{-1}$.
Then, with taking $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}n^{-1}$, (3.1) holds by the fact that $1-(1-\alpha)H>\frac{1}{2}.$
Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). Similarly to (4.7), we can write
${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=$ $:I_{1}+I_{2}.$
The two terms are treated in the exactly same manner, so we only detail $I_{1}$. Similarly to (4.8), we get
$\|I_{1}\|_{L_{\omega}^{p}}\leqslant$ $N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}$
$\qquad+(r-s_{3})^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}\mathrm{d}r.$ (4.22)
Similar to (4.10), we have
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|\|_{L^{p}_{\omega}}.$ (4.23)
We note that
$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}$ $=\int_{k_{n}(r)}^{r}b(B^{H}_{k_{n}(t)}+\varphi_{k_{n}(t)}^{n})\mathrm{d}t=(r-k_{n}(r))b(B_{k_{n}(r)}^{H}+\varphi^{n}_{k_{n}(r)})\in\mathcal{F}_{k_{n}(r)}.$
When $s_{1}\in[k_{n}(r),r],$ we have
$\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})=0;$
[…]
${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
$=$ $(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}b(B_{k_{n}(r)}^{H}+\varphi_{k_{n}(r)}^{n})|$
$\leqslant$ $2(r-k_{n}(r)){\mathbb{E}}^{s_{1}}|b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})-b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+\varphi_{s_{1}}^{n})|$
$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}{\mathbb{E}}^{s_{1}}(|B_{k_{n}(r)}^{H}-{\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-\varphi_{s_{1}}^{n}|^{\alpha}).$
Moreover, using (3.5) and (4.2), we have
${\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|$
$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}(|k_{n}(r)-s_{1}|^{\alpha H}+|k_{n}(r)-s_{1}|^{\alpha})$
$\leqslant$ $\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}$
where we used the fact $H>1$ in the second inequality. Plugging it into (4.23), we get
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}.$ (4.24)
Meanwhile (4.9) and (4.2) yield
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\leqslant N|k_{n}(r)-s_{1}|^{1+\alpha H}\leqslant N|r-s_{1}|^{1+\alpha H},$ (4.25)
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})\|_{L_{\omega}^{p}}\leqslant N\|\varphi_{\cdot}^{n}-\varphi_{k_{n}(\cdot)}^{n}\|_{C_{p}^{0}}\leqslant N\frac{\|b\|_{C^{0}}}{n}.$ (4.26)
Applying (4.24), (4.25) and (4.26) into (4.22) gives us
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}(r-s_{1})^{\alpha}\mathrm{d}r+(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r$
$\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)}.$ (4.27)
With the same bound on $I_{2}$ and Assumption˜2.1 implying $(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)>1$, (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
Lastly we claim that the process $\mathcal{A}$ in (3.3) is indeed given by
$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-b(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.28)
To prove this claim, it suffices to show (3.4). Similarly to (4), we have
$\mathcal{A}t-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{r})\mathrm{d}r$
$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
Then the estimates can be done similarly to that in Lemma˜4.2.
Therefore, the claim holds by the uniqueness from Lemma˜3.1. Finally, by Lemma˜3.1, we can conclude (4.18). ∎
[…]
Lemma 4.4.
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$, we have
$\left\|\mathcal{E}^{b,n,3}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.29)
with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
Proof.
In order to apply Lemma˜3.1, this time we set
$A_{s,t}={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})-b(B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
When $|t-s|\leqslant\frac{10}{n}$, by (3.6), we have for any $\epsilon>0$
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}\big\||B^{H}_{r}-B^{H}_{k_{n}(r)}|^{\alpha}\big\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|\cdot\frac{1}{n^{\alpha}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}\frac{1}{n^{\alpha+\frac{1}{2}-\varepsilon}}.$
For $H>1$ and $\alpha>1-\frac{1}{2H}$ we can take such sufficiently small $\varepsilon>0$ so that $\epsilon\in(0,\alpha+\frac{1}{2H}-1)$, which implies
$\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}.$ (4.30)
When $|t-s|>\frac{10}{n}$, we have
$A_{s,t}$ $=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad+\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=:IV_{1}+IV_{2}.$
For $IV_{1}$, by (3.9), (2.1) and (3.6), we have
$\|IV_{1}\|_{L^{p}_{\omega}}$ $\leqslant\int_{s_{4}}^{s_{6}}\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{1}}\|{\mathbb{E}}^{s_{1}}(B_{r}^{H}-B_{k_{n}(r)}^{H})\|_{L^{p}_{\omega}}\mathrm{d}r$
$\leqslant\frac{N\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{6}}(r-s_{1})^{-(1-\alpha)H}\mathrm{d}r$
$\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H}.$ (4.31)
For $IV_{2}$, (3.10) with $\delta=1$ gives us
$\|\mathcal{P}^{H}_{t}b-\mathcal{P}_{s}^{H}b\|_{{\mathcal{C}}^{0}}\leqslant Ns^{-(2-\alpha)H}|t^{2H}-s^{2H}|\|b\|_{{\mathcal{C}}^{\alpha}},$
it implies
$|IV_{2}|$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(k_{n}(r)-s_{1})^{-(2-\alpha)H}\left((r-s_{1})^{2H}-(k_{n}(r)-s_{1})^{2H}\right)\mathrm{d}r.$
Moreover, by $k_{n}(r)-s_{1}\asymp t-s,$ and
$\big||r-s_{1}|^{2H}-|k_{n}(r)-s_{1}|^{2H}\big|\leqslant N|r-k_{n}(r)||r-s_{1}|^{2H-1}\leqslant\frac{N}{n}|t-s|^{2H-1},$
we have
$|IV_{2}|\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(t-s)^{2H-1-(2-\alpha)H}\mathrm{d}r=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\alpha H}.$ (4.32)
Then (4.30) togehter with (4.31) and (4.32) verifies (3.1) of Lemma˜3.1 with taking $C_{1}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$, since $((1-(1-\alpha)H)\wedge(\alpha H\big))>\frac{1}{2}$.
Next we verify (3.2). Let $(s,u,t)\in\overline{[0,1]}_{1}^{3}$. Recall the definition of $s_{i},i=1,\dots,6$ in (4.3). Similarly to (4.7), we have
[…]
$\qquad-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
[…]
$\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{r}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{2}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
$=:V_{1}+V_{2}.$
Again we observe that the above two terms can be treated in the exactly same way, so we only detail $V_{1}$.
Applying (3.3) with taking
$x_{1}={\mathbb{E}}^{s_{3}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)},x_{2}={\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n},$
$x_{3}={\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)},x_{4}={\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n},$
we get
$\|V_{1}\|_{L^{p}_{\omega}}$
$\leqslant N\int_{s_{4}}^{s_{5}}\|b\|_{{\mathcal{C}}^{\alpha}}(r-s_{3})^{-H(2-\alpha)}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}\mathrm{d}r.$ (4.33)
By Cauchy-Schwarz inequality,
${\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]$
$\leqslant({\mathbb{E}}^{s_{1}}({\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)}))^{2})^{\frac{1}{2}}({\mathbb{E}}^{s_{1}}({\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})^{2})^{\frac{1}{2}}.$
By Jensen inequality, we obtain
$({\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)}))^{2}\leqslant$ ${\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})^{2},$
$({\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})^{2}=$ $({\mathbb{E}}^{s_{3}}(\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}))^{2}\leqslant{\mathbb{E}}^{s_{3}}(\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})^{2}.$
Therefore, by (4.2), we have
${\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]$
$\leqslant({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot({\mathbb{E}}^{s_{1}}(\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})^{2})^{\frac{1}{2}}$
$\leqslant N({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot(k_{n}(r)-s_{1})^{1+\alpha H}.$
For $p\geqslant 2$, (2.1) and (3.6) imply
$\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B^{H}_{k_{n}(r)})||{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}$ (4.34)
$\leqslant N(r-s_{1})^{1+\alpha H}\left\|\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L_{\omega}^{2}|\mathcal{F}_{s_{1}}}\right\|_{L^{p}_{\omega}}$
$=N(r-s_{1})^{1+\alpha H}\|{\mathbb{E}}^{s_{1}}|B_{r}^{H}-B_{k_{n}(r)}^{H}|^{2}\|^{\frac{1}{2}}_{L_{\omega}^{\frac{p}{2}}}$
$\leqslant N(r-s_{1})^{1+\alpha H}\|B_{r}^{H}-B^{H}_{k_{n}(r)}\|_{L^{p}_{\omega}}$
$\leqslant\frac{N}{n}(r-s_{1})^{1+\alpha H}.$
Therefore, plugging (4.34) into (4), we have
$\|V_{1}\|_{L^{p}_{\omega}}$ $\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{2+(2\alpha-2)H}.$
The same bound holds on $V_{2}$.
Since Assumption˜2.1 implies $2+(2\alpha-2)H>1$, we can obtain that (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
In order to show that the process $\mathcal{A}$ in (3.3) is given by
$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi^{n}_{k_{n}(r)})-b(B^{H}_{k_{n}(r)}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r,$
it suffices to show (3.4). For doing so, we first write
$\mathcal{A}_{t}-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B_{r}^{H}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}_{r-s_{1}}^{H}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
$\quad-\int_{s}^{t}b(B^{H}_{k_{n}(r)}+\varphi^{n}_{k_{n}(r)})-\mathcal{P}^{H}_{(k_{n}(r)-s_{1})\vee 0}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r.$
Then similar to what we have done in Lemma˜4.2 (that is, (4), (4.16) and (4.17)), we can conclude the claim by the uniqueness from Lemma˜3.1.
In the end all of the conditions from Lemma˜3.1 are verified, which proves the desired result. ∎
With Lemma˜4.2, Lemma˜4.3 and Lemma˜4.4 at hand we are ready to give:
Proof of Theorem˜2.2.
By (4.4), (4.18) and (4.29), we see that
$\|(X-X^{n})_{t}-(X-X^{n})_{s}\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon},$
which implies
$[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}\leqslant N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
Therefore, we have
$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}$ $\leqslant|(X-X^{n})_{S}|+2[X-X^{n}]_{C_{p}^{\frac{1}{2}}[S,T]}$
$\leqslant|(X-X^{n})_{S}|+N(\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[S,T]}+n^{-1})(T-S)^{\varepsilon}.$
Fix $T-S=\Delta$ small enough and we obtain
$\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}[S,T]}\leqslant N(|(X-X^{n})_{S}|+n^{-1}).$
Dividing $[0,1]$ into $0,\Delta],[\Delta,2\Delta],\ldots$, yields that
$\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}[0,1]}\leqslant N(|x_{0}-x_{0}^{n}|+n^{-1}).$
5. Optimality
In the following we present a result concerning the optimal convergence rate of EM (1.9) for rather smooth $b\in{\mathcal{C}}^{1}$.
Theorem 5.1.
Suppose $b\in{\mathcal{C}}^{1}$ and $x_{0}^{n}=x_{0}$. Then for any $t\in[0,1]$, a.s.
$\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})=:c(t)$ (5.1)
exists and $c(t)$ satisfies
$c^{\prime}(t)=\nabla b(X_{t})c(t)+\frac{1}{2}\nabla b(X_{t})(b(X_{t})+(B_{t}^{H})^{\prime}),\quad c(0)=0.$
Proof.
By (1.3) and (1.9), we have
$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}n(b\left(X_{r}\right)-b\left(X_{r}^{n}\right))+n(b\left(X_{r}^{n}\right)-b(X_{k_{n}(r)}^{n}))\mathrm{d}r$
$=\int_{0}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))n(X_{r}-X_{r}^{n})$
$\qquad\quad+\nabla b(X_{k_{n}(r)}^{n}+\theta(X_{r}^{n}-X_{k_{n}(r)}^{n}))n(X_{r}^{n}-X_{k_{n}(r)}^{n})\mathrm{d}\theta\mathrm{d}r.$
By fundamental theory of calculus we can write
$n(X_{t}-X_{t}^{n})$ $=\int_{0}^{t}\exp\left(\int_{s}^{t}\int_{0}^{1}\nabla b(X_{r}^{n}+\theta(X_{r}-X_{r}^{n}))d\theta dr\right)$
$\qquad\int_{0}^{1}\nabla b(X_{k_{n}(s)}^{n}+\theta(X_{s}^{n}-X_{k_{n}(s)}^{n}))n(X_{s}^{n}-X_{k_{n}(s)}^{n})\mathrm{d}\theta\mathrm{d}s=:c^{n}(t).$
Following from (1.9) and the fact a.s. $B^{H}\in{\mathcal{C}}^{H-\varepsilon}([0,1])$ for sufficiently small $\varepsilon>0$ by Kolmogorov continuity criteria, which means there exists a small enough $\epsilon^{\prime}>0$ so that a.s. $B_{H}\in{\mathcal{C}}^{1+\epsilon^{\prime}}$, we have a.s. (denote $\{c\}:=c-\lfloor c\rfloor$ for $c\in{\mathbb{R}}_{+}$)
$n(X_{s}^{n}-X_{k_{n}(s)}^{n})$ $=n(s-k_{n}(s))b(X_{k_{n}(s)}^{n})+n(B_{s}^{H}-B_{k_{n}(s)}^{H})$
$=\{ns\}b(X_{k_{n}(s)}^{n})+\{ns\}(B_{s}^{H})^{\prime}+o(n^{-\varepsilon^{\prime}});$
together with the condition that $\nabla b$ is continuous and bounded, moreover $X_{r}^{n}\rightarrow X_{r}$ and $X_{k_{n}(s)}^{n}\rightarrow X_{s}$ a.s. from Theorem˜2.2, dominated convergence theorem shows as $n\rightarrow\infty$, we have
$c^{n}(t)$ $-\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\rightarrow 0\quad a.s.$ (5.2)
Besides, observe
$\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s$
$\rightarrow\frac{1}{2}\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\quad a.s.$
since $\{n\cdot\}$ converges to $\frac{1}{2}$ weakly in $L^{2}([0,t])$ as $n\rightarrow\infty$, the above combining with (5.2) completes the proof. ∎
Remark 5.2.
Notice from Theorem˜5.1, we verify that indeed order $n$ reaches the optimal rate of EM (1.9), although for $b\in{\mathcal{C}}^{1}$ instead of assuming $b\in{\mathcal{C}}^{\alpha}$ for $\alpha<1$. Putting Theorem˜2.2 and Theorem˜5.1 together we can evidently say that order $n$ reaches the optimal (upper bound) rate of EM also for $b\in{\mathcal{C}}^{\alpha}$ simply by the embedding ${\mathcal{C}}^{1}\subset{\mathcal{C}}^{\alpha}$. However, concerning the lower bound of (1.9) for $\alpha<1$ we can not say too much so far.
Acknowledgment
CL is supported by Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 563883019. KS is grateful to the financial supports by National Key R & D Program of China (No. 2022YFA1006300) and the financial supports of the NSFC (No. 12426205, No. 12271030).
We are grateful to Prof. Rongchan Zhu (Beijing Institute of Technology) for her valuable suggestions.